OpenStudy (rubylovesjaune):

1. Bror simplifies the expression x^2 + 3x +2 / x+2 as follows: x^2 + 3 + 1 / 1+2 = x^2 + 4 / 1 = x^2 +2 His teacher tells him that his answer is incorrect. State what Bror did wrong, and then simplify the expression correctly. How do you do this problem?

OpenStudy (hyuna301):

(x^2 + 3x + 2)/(x+2) = (x^2 + x + 2x + 2)/(x+2) = [(x+2) (x+1)]/(x+2) = x +1 ,but x should not equal to 2 because the eq. will become undefined

OpenStudy (anonymous):

@Hyuna301 looke please my opinion that for x= -2 will be thie equation undefined bc. than will get the denominator equal zero - please correct your answer

OpenStudy (hyuna301):

oh.. i just missed it by mistake ..sorry

OpenStudy (radar):

The problem states it is an expression, in other words there was not originally an equal sign. All Bror was requested to do, was simplify the expression, which he did incorrectly. You on the other hand got the correct results.

OpenStudy (mathstudent55):

Bror did this: \(\dfrac{x^2 + 3x +2}{x+2} \) \(=\dfrac{x^2 + 3 + 1}{1+2} \) You can't just divide 3x/x = 1, and I don't know what was done to the 2 in the numerator and the 2 in the denominator. \(=\dfrac{x^2 + 4}{1} \) In this step 3 + 1 = 4 in the numerator. How does 1 + 2 = 1 in the denominator? Of course, none of it is correct because the previous step was wrong. \(=x^2 +2 \) How can dividing something by 1 change it. How can dividing x^2 + 4 by 1 give x^2 + 2? Every step of Bror's answer is incorrect because a new mistake is made in every step in addition to the mistakes made in the previous steps.

OpenStudy (mathstudent55):

The correct solution: \(\dfrac{x^2 + 3x +2}{x+2} =\) Factor the numerator: \(=\dfrac{(x + 2)(x + 1)}{x+2} \) Now you can divide the numerator and denominator by the factor x + 2: \(= x + 1\)

OpenStudy (radar):

Agree, Bror messed up big time. Avoid having Bror help you with your homework.